OpenStudy (anonymous):

Consider the quadratic equation: y=−2x 2 −32x+2 . Write the equation in the form y=a(x−h) 2 +k by completing the square.

OpenStudy (campbell_st):

well take out -2 as a common factor \[y = -2(x^2 + 16x ) + 2\] complete the square on the quadratic \[y = -2(x^2 +16x + (\frac{16}{2})^2) + 2 - (\frac{16}{2})^2\] hope this helps.

OpenStudy (anonymous):

yes it does I actually got an answer but its telling me im wrong can i tell u what i got. >

Directrix (directrix):

@spimentel Post your work here, please, so we can take a look at it.